What is the fluid speed in a fire hose with a 9.05-cm diameter carrying 55 L of water per second? 

Round your answer to 3 decimal places


1
Expert's answer
2021-04-23T11:06:28-0400

"Q=55\space l/s=55\times10^{-3}\space m^3/s"

"d=9.05\space cm=9.05\times10^{-2}\space m"

"Q=Av"

"A:" Area of cross-section of pipe

"v:" velocity of flow

"v=\dfrac{Q}{A}=\dfrac{55\times10^{-3}}{\dfrac{\pi d^2}{4}}"

"v=8.550\space m/s"